OpenStudy (anonymous):

Transform this equation to slope-intercept form: 3x + 9y = 18

OpenStudy (ahsome):

Slope intercept form is basicly \(y\) is on one side by itself, and \(x\) is on the other side Equation: \(3x + 9y = 18 \) Move the \(3x\) to the other side \[3x + 9y = 18\]\[9y=18-3x\]Now, we normally put the \(x\) valye infront, so simply move it infront\[9y=18-3x\]\[9y=-3x+18\]Tada :)

OpenStudy (anonymous):

y=mx+b so subract 3x from both sides of the equation, then divide everything by nine